Please solve, thanks.Growth Companys current share price is $20.15 and it is expected to pay a $0.95 dividend per share next year. After that, th
Answer #1

1.
=0.95/20.15+4.1%
=8.81464%

2.
=2.20/28.20
=7.80142%

3.
=6.9%

4.
=5.2*20.15+1.4*28.20+19.8
=164.06 million

5.
=(19.8*(6.9%*(1-40%))+1.4*28.20*(2.20/28.20)+5.2*20.15*(0.95/20.15+4.1%))/(5.2*20.15+1.4*28.20+19.8)
=8.00664%